OpenStudy (anonymous):

Given the function f(x) = -x^2-6x-2 i. Determine the vertex; ii. List the axis of symmetry; iii. Calculate the y-intercept; iv. Use the axis of symmetry and the y-intercept to find an additional point on the graph; v. Graph the function Please show all of your work. Given the function f(x) = -x^2-6x-2 i. Determine the vertex; ii. List the axis of symmetry; iii. Calculate the y-intercept; iv. Use the axis of symmetry and the y-intercept to find an additional point on the graph; v. Graph the function Please show all of your work. @Mathematics

OpenStudy (anonymous):

f(x) = -x^2-6x-2 the negative sign in x might tells us that the parabola opens downward vertex x=-b/2a x=-(-6)/2(-1)=6/-2=-3 sub this to f(x) gives f(3)= -(-3)^2-6(-3)-2=-9+18-2=7 therefore vertex(h,k)=(-3,7) ans 110 axis of sym is at y

OpenStudy (anonymous):

11) axis of sym is at y

OpenStudy (anonymous):

f(x) = -x^2-6x-2 when x =0, y=f(0)=-2 ,,,y intercept

OpenStudy (anonymous):

y=f(x) = -x^2-6x-2 when y=0, we can solve for the x intercepts, by using the formula x=[-b-+sqrt(b^2 -4ac)]/2a a=-1,b=-6 and c=-2 we get x=-0.354,-5.6

OpenStudy (anonymous):

to graph this use these points x=-0.354,-5.6, x =0, y=f(0)=-2 ,,,y intercept

OpenStudy (anonymous):

also use the points of the vertex x=-3, y=7

OpenStudy (anonymous):